# Factoriality of groupoid von Neumann algebras

### v2

@@ 1 Introduction
Discrete measured groupoids have historically played an important role in the study of von Neumann algebras. In [Mac63], Mackey introduced the notion of ergodic groupoids. An early breakthrough in this direction was due to Feldman and Moore who showed in [FM77] that every principal discrete measured groupoid, i.e., countable equivalence relation arises as the orbit equivalence relation of a countable group action. In [FM77a] they defined the von Neumann algebra $L(\mathcal{R})$ associated to such an equivalence relation $\mathcal{R}$ on a standard probability space $(X,\mu)$. They showed that the abelian subalgebra $L^{\infty}(X,\mu)$ is a so-called Cartan subalgebra of $L(\mathcal{R})$. In-fact they showed that any von Neumann algebra with a Cartan subalgebra arises from such an equivalence relation together with a 2-cocycle, thus giving a complete characterization.
@@
In this article, we study the question: when is the von Neumann algebra associated to a discrete measured groupoid a factor? We give a characterization of factoriality purely in terms of the groupoid. Up until this point, such a characterization was only known in special cases. For a countable group $\Gamma$, it is well known that $L(\Gamma)$ is a factor if and only if the group has infinite conjugacy classes (such groups are generally called icc). In the case of a countable equivalence relation $\mathcal{R}$ it was proved in [FM77a] that the factoriality of the von Neumann algebra $L(\mathcal{R})$ is equivalent to the ergodicity of $\mathcal{R}$. Finally, when $\Gamma\curvearrowright(X,\mu)$ is an ergodic (not necessarily free) action of a countable group on a standard probability space, a characterization of the factoriality of the crossed product $L^{\infty}(X,\mu)\rtimes\Gamma$ was given by Vaes in [Vae20].
The organisation of the first part of the paper is as follows. We state some preliminaries on discrete measured groupoids and their von Neumann algebras in Section 2, and then gather some folklore results on bases of groupoids in Section 3. A lot of these results have already been used intrinsically in the literature, for example in [Ana13]. In Section 4, we define the notion of conjugacy classes for Borel subsets of the isotropy subgroupoid. For such a Borel subset $E$ we show that the set $\Omega_{E}$ consisting of all conjugates of $E$ is a Borel subset and we call such a conjugacy class finite if it has finite measure. In Definition 4.2 we define groupoids satisfying the infinite conjugacy class (icc) condition if any such subset with a finite conjugacy class is essentially a subset of the unit space. Then we prove the following theorem, which is the main result of this paper.
@@
[…]
Similarly for factoriality, beyond the principal case the 2-cocycle creates a non-trivial obstruction. A characterization of factoriality for twisted group von Neumann algebras was given in [Kle62]. In Section 5.3, we introduce the notion of central subsets of the isotropy subgroupoid. If $\omega$ is a 2-cocycle on the groupoid, we call such a conjugation invariant Borel subset $E\subset\operatorname{Iso}(\mathcal{G})$ central if there exists a function $f:E\rightarrow\mathbb{C}$ such that:
$f(ghg^{-1})=\overline{\omega(ghg^{-1},g)}\omega(g,h)f(h)$
for all $h\in E$ and all elements $g$ with the same source as $h$. In Theorem 5.5 and Corollary 5.6, we give a characterization of factoriality for cocycle twisted groupoid von Neumann algebras in terms of these central subsets. We also introduce a similar condition as in [Kle62] for twisted discrete measure groupoids, and show that it is a necessary condition for the factoriality of the twisted groupoid von Neumann algebra.
@@
Acknowledgements: The authors would like to thank Stefaan Vaes for his valuable comments on the article.
@@ 2 Preliminaries
  • $\mathcal{G}^{(0)}$ is a Borel subset of $\mathcal{G}$.
  • the source and target maps $s$ and $t$ are countable-to-one Borel maps,
  • composition and inversion are Borel maps.
A (non-singular) discrete measured groupoid $(\mathcal{G},\mu)$ is a discrete Borel groupoid $\mathcal{G}$ together with a probability measure $\mu$ on $\mathcal{G}^{(0)}$ such that the following two $\sigma$-finite measures on $\mathcal{G}$ are equivalent:
[…]
@@
Let $(\mathcal{G},\mu)$ be a discrete measured groupoid such that $(\mathcal{G}^{(0)},\mu)$ is a finite measure space. To construct the groupoid von Neumann algebra, we start by defining partial isometries $\lambda_{A}\in B(L^{2}(\mathcal{G},\mu))$ for each bisection $A\subset\mathcal{G}$ by
$(\lambda_{A}1_{E})(g)=1_{A}*1_{E}=1_{A\cdot E}$ (1)
for every subset $E\subset\mathcal{G}$ with $\mu_{s}(E)<\infty$. The groupoid von Neumann algebra is defined by $L(\mathcal{G}):=\{\lambda_{A}\ |\ A\subset\mathcal{G}\text{ Borel bisection}\}^{\prime\prime}$. Whenever $A$ and $B$ are bisections, one sees that $\lambda_{A}\lambda_{B}=\lambda_{A\cdot B}$. Note that whenever $A\subset\mathcal{G}$ is a Borel subset with $\mu_{s}(A)<\infty$, it defines an element of $L(\mathcal{G})$ which we also denote by $\lambda_{A}$ as it is defined by the formula (1).
@@
Note that $L(\mathcal{G})$ includes a copy of $L^{\infty}(\mathcal{G}^{(0)})$ as a regular subalgebra. Whenever $A\subset\mathcal{G}$ is a Borel bisection and $F\in L^{\infty}(\mathcal{G}^{(0)},\mu)$, one has the relation $\lambda_{A}F\lambda_{A}^{*}=(F\circ\sigma_{A}^{-1})\cdot 1_{t(A)}$. Moreover, there is always a faithful normal conditional expectation $E:L(\mathcal{G})\to L^{\infty}(\mathcal{G}^{(0)},\mu)$ given by $a\mapsto a1_{\mathcal{G}^{(0)}}|_{\mathcal{G}^{(0)}}$. The fact that $E(a)\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ follows from Lemma 3.4.
@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
The following proposition is a consequence a theorem of Lusin-Novikov, see e.g. [Kec95, Theorem 18.10], as noted by Anatharaman-Delaroche in [Ana13]. We give a proof for convenience.
@@
Since $s:\mathcal{G}\to\mathcal{G}^{(0)}$ is countable-to-one, we can use the Lusin-Novikov Theorem [Kec95, Theorem 18.10] to find a countable collection $(U_{n})_{n}$ of Borel sets $U_{n}\subset\mathcal{G}$ such that $s:U_{n}\to s(U_{n})$ is a Borel isomorphism for every $n$ and such that $\mathcal{G}=\bigsqcup_{n}U_{n}$. Indeed, setting $P:=\{(x,y)\in\mathcal{G}^{(0)}\times\mathcal{G}\ |\ x=s(y)\}$, by Lusin-Novikov there is a partition $P=\bigsqcup_{n}P_{n}$. Denoting by $\pi$ the projection from $\mathcal{G}_{0}\times\mathcal{G}$ onto $\mathcal{G}$, set $U_{n}=\pi(P_{n})$. Similarly, there exists a countable collection $(V_{n})_{n}$ of Borel subsets $V_{n}\subset\mathcal{G}$ such that $t:V_{n}\to t(V_{n})$ is a Borel isomorphism for every $n$ and such that $\mathcal{G}=\bigsqcup_{n}V_{n}$. Define =$\mathcal{B}^{\prime}=\{(U_{n}\cap V_{m})\setminus\mathcal{G}_{0}\}_{n,m}$ and set =$\mathcal{B}=\{A\cap B^{-1}\ |\ A,B\in\mathcal{B}^{\prime}\}\cup\{\mathcal{G}^{(0)}\}$. ∎
@@
where $f\cdot 1_{B}$ denotes pointwise multiplication and convergence of the series is in $L^{2}$-norm.
@@
$\langle 1_{B},1_{C}\rangle=\langle\lambda_{B}1_{\mathcal{G}^{(0)}},\lambda_{C}1_{\mathcal{G}^{(0)}}\rangle=\mu\left(\left(B\cdot C^{-1}\right)\cap\mathcal{G}^{(0)}\right)$
and $\left(B\cdot C^{-1}\right)\cap\mathcal{G}^{(0)}=\emptyset$ whenever $B\neq C$. The decomposition of $f\in L^{2}(\mathcal{G},\mu_{s})$ follows from the fact that $(\mathcal{G},\mu_{s})=\left(\bigsqcup_{B\in\mathcal{B}}B,\sum_{B\in\mathcal{B}}\mu_{s}|_{B}\right)$. Finally, a computation shows that for any $f\in\mathcal{B}$ it holds that
@@
$j:L(\mathcal{G})\to L^{2}(\mathcal{G},\mu):a\mapsto j(a):=a1_{\mathcal{G}^{(0)}}.$ (3)
It is not difficult to check that $j$ is a continuous linear map. To see that $j$ is injective, denote by $E$ the conditional expectation from $L(\mathcal{G})$ onto $L^{\infty}(\mathcal{G}^{(0)},\mu)$, define the state $\phi_{\mu}$ on $A(\mathcal{G},\mu)$ by setting
[…]
@@
Now suppose $a\in L(\mathcal{G})$ with $\|a\|\leq 1$. Let $(a_{n})_{n}$ be a sequence in $A(\mathcal{G},\mu)$ be a sequence with $\|a_{n}\|\leq 1$ for all $n$ and such that $a_{n}\to a$ in the SOT. By definition, we get that $a_{n}1_{\mathcal{G}^{(0)}}\to a1_{\mathcal{G}^{(0)}}$ in $L^{2}$-norm. Then also $a_{n}1_{\mathcal{G}^{(0)}}s|_{B}^{-1}\to a1_{\mathcal{G}^{(0)}}a|_{B}^{-1}$ in $L^{2}$-norm. Now it suffices to note that from $\|a\|\leq 1$ it follows that $\|a_{n,B}\|_{\infty}=\|a_{n}1_{\mathcal{G}^{(0)}}s|_{B}^{-1}\|_{\infty}\leq 1$. Then, since the $\|\cdot\|_{2}$ and the $\sigma$-strong* topology agree on bounded sets, we get that $a_{n,B}\to a_{B}$ in the SOT, which implies that $a_{B}\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ since all $a_{n,B}\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ by the first part of the proof.
@@
We claim that the net $a_{\mathcal{F}}$ converges to $a$ in the sharp norm. We first show that $\phi((a-a_{\mathcal{F}})^{*}(a-a_{\mathcal{F}}))\to 0$ as $\mathcal{F}$ tends to infinity. This follows from noting that
$\phi((a-a_{\mathcal{F}})^{*}(a-a_{\mathcal{F}}))=\langle j(a-a_{\mathcal{F}}),j(a-a_{\mathcal{F}})\rangle$ $=\left\|\sum_{B\in\mathcal{B}\setminus\mathcal{F}}a^{B}\ast 1_{B}\right\|_{2}$
To see that $\|E((a-a_{\mathcal{F}})(a-a_{\mathcal{F}})^{*}))\|_{2}\to 0$ as $\mathcal{F}$ tends to infinity, note that
@@
so it suffice to check that $(a_{\mathcal{F}})^{*}$ gives the appropriate Fourier decomposition of $a^{*}$. To this end, a simple calculation shows that
@@
Finally, to see that the Fourier coefficients satisfy $E(a\lambda_{B}^{*})=a^{B}$, fix a $B\in\mathcal{B}$ and observe that whenever $\mathcal{F}\subset\mathcal{B}$ is a finite set containing $B$, it holds that $E(a_{\mathcal{F}}\lambda_{B}^{*})=a^{B}$. The $E(\cdot\lambda_{B}^{*})$ is a normal map, hence $\|\cdot\|^{\sharp}$-continuous on bounded subsets. Since the net $(a_{\mathcal{F}})$ is bounded in sharp norm by $\|a\|^{\sharp}$, we indeed get that $(a\lambda_{B}^{*})=\lim_{\mathcal{F}}E(a_{\mathcal{F}}\lambda_{B}^{*})=a^{B}$. ∎
@@
Let $\mathcal{G}$ be a discrete Borel groupoid and let $\mathcal{B}$ be a basis for $\mathcal{G}$. Suppose $C\subset\mathcal{B}$ is a bisection. The set of conjugates $C\mathcal{B}C^{-1}=\{CBC^{-1}\ |\ B\in\mathcal{B}\}$ also forms a basis for $C\mathcal{G}C^{-1}$. If $\mathcal{B}$ is symmetric, then so is $C\mathcal{B}C^{-1}$.
@@
First let us show that elements of $C\mathcal{B}C^{-1}$ are mutually disjoint. Let $B,D\in\mathcal{B}$. Suppose that $a,b,c,d\in C$ and $x\in B$ and $y\in D$ such that $axb^{-1}=cyd^{-1}$. Note in particular that $t(a)=t(c)$ and that $t(b)=t(d)$. However, each $a,b,c,d$ belong to the same bisection. By injectivity of the target map, we get that $a=c$ and $b=d$. Thus, $x=s(c)ys(d)=y\in B\cap D$. Therefore, $B=D$. Finally, observe that
@@
We argue by contradiction. Suppose $j(a)$ is not supported on $\operatorname{Iso}(\mathcal{G})$. Then we find a Borel subset $E\subset\mathcal{G}\setminus\operatorname{Iso}(\mathcal{G})$ such that $j(a)(g)\neq 0$ for all $g\in E$. Since $\mathcal{B}$ is countable and mutually disjoint, there is a $B_{0}\in\mathcal{B}$ such that $\mu_{s}(E\cap B_{0})>0$. Note that for every $x\in s(E\cap B_{0})$ we have that $a_{B_{0}}(x)\neq 0$. By the centrality of $a$, we get for any Borel $F\subset s(E\cap B_{0})$ that
$j(\lambda_{F}a)=\sum_{B\in\mathcal{B}}1_{F\cdot B}\ast a_{B}=\sum_{B}1_{B}\ast(1_{F}\circ\sigma_{B})a_{B}=j(a\lambda_{F})=\sum_{B\in\mathcal{B}}1_{B}\ast a_{B}1_{F}.$
Since $a_{B}1_{F}$ and $(1_{F}\circ\sigma_{B})a_{B}$ are both elements of $L^{\infty}(\mathcal{G}^{(0)},\mu)$, looking at the restriction of $j(\lambda_{F}a)$ on each basis element gives us the identity $a_{B}1_{F}=(1_{F}\circ\sigma_{B})a_{B}$ for all $B\in\mathcal{B}$. Hence in particular $1_{F}=(1_{F}\circ\sigma_{B_{0}})$, meaning that for any $x\in s(E\cap B_{0})$, $x\in F$ if and only if $\sigma_{B_{0}}(x)\in F$. Since this holds for any Borel subset $F\subset s(E\cap B)$, we get that $x=t\circ s|_{B_{0}}^{-1}(x)$ for all $x\in s(E\cap B)$. In other words, $E\cap B\subset\operatorname{Iso}(\mathcal{G})$, which is a contradiction. ∎

### v3

@@ 1 Introduction
Discrete measured groupoids have historically played an important role in the study of von Neumann algebras. In [Mac63], Mackey introduced the notion of ergodic groupoids. An early breakthrough in this direction was due to Feldman and Moore who showed in [FM77a] that every principal discrete measured groupoid, i.e. countable equivalence relation, arises as the orbit equivalence relation of a countable group action. In [FM77b] they defined the von Neumann algebra $L(\mathcal{R})$ associated to such an equivalence relation $\mathcal{R}$ on a standard probability space $(X,\mu)$. They showed that the abelian subalgebra $L^{\infty}(X,\mu)$ is a so-called Cartan subalgebra of $L(\mathcal{R})$. In fact they showed that any von Neumann algebra with a Cartan subalgebra arises from such an equivalence relation together with a 2-cocycle, thus giving a complete characterization.
@@
In this article, we study the question: when is the von Neumann algebra associated to a discrete measured groupoid a factor? We give a characterization of factoriality purely in terms of the groupoid. Up until this point, such a characterization was only known in special cases. For a countable group $\Gamma$, it is well known that $L(\Gamma)$ is a factor if and only if the group has infinite conjugacy classes (such groups are generally called icc). In the case of a countable equivalence relation $\mathcal{R}$ it was proved in [FM77b] that the factoriality of the von Neumann algebra $L(\mathcal{R})$ is equivalent to the ergodicity of $\mathcal{R}$. Finally, when $\Gamma\curvearrowright(X,\mu)$ is an ergodic (not necessarily free) action of a countable group on a standard probability space, a characterization of the factoriality of the crossed product $L^{\infty}(X,\mu)\rtimes\Gamma$ was given by Vaes in [Vae20].
The organisation of the first part of the paper is as follows. We state some preliminaries on discrete measured groupoids and their von Neumann algebras in Section 2, and then gather some folklore results on bases of groupoids in Section 3. A lot of these results have already been used intrinsically in the literature, for example in [AD13]. In Section 4, we define the notion of conjugacy classes for Borel subsets of the isotropy subgroupoid, which consists of all elements whose source and target are equal. For such a Borel subset $E$ we show that the set $\Omega_{E}$ consisting of all conjugates of $E$ is a Borel subset and we call such a conjugacy class finite if it has finite measure. In Definition 4.2 we define groupoids satisfying the infinite conjugacy class (icc) condition if any such subset with a finite conjugacy class is essentially a subset of the unit space. Then we prove the following theorem, which is the main result of this paper.
@@
The situation is more complicated when a non-trivial cocycle twist is involved. Indeed for a 2-cocycle $\omega$ on $\mathcal{G}$, one can define the twisted groupoid von Neumann algebra as we do in Section 5. However, a lot of statements for such twisted von Neumann algebras become more challenging become more challenging to prove. For example, as observed in [Hah78], the result in [Con75] does not address the question: ‘Can every factor can be realized as a cocycle twisted groupoid von Neumann algebra?’ which remained open for a long time. In fact, only recently in [DV24], the third named author and Stefaan Vaes constructed a family of II1 factors that are not virtually isomorphic to any twisted groupoid von Neumann algebra, even allowing arbitrary amplifications.
Similarly for factoriality, beyond the principal case the 2-cocycle creates a non-trivial obstruction. A characterization of factoriality for twisted group von Neumann algebras was given in [Kle62]. In Section 5.3, we introduce the notion of central subsets of the isotropy subgroupoid. If $\omega$ is a 2-cocycle on the groupoid, we call such a conjugation invariant Borel subset $E\subset\operatorname{Iso}(\mathcal{G})$ central if there exists a function $f:E\rightarrow\mathbb{C}$ such that $f(ghg^{-1})=\overline{\omega(ghg^{-1},g)}\omega(g,h)f(h)$ for elements $h\in E$ and all elements $g$ with the same source as $h$. In Theorem 5.5 and Corollary 5.6, we give a characterization of factoriality for cocycle twisted groupoid von Neumann algebras in terms of these central subsets. We also introduce a similar condition as in [Kle62] for twisted discrete measure groupoids, and show that it is a necessary condition for the factoriality of the twisted groupoid von Neumann algebra.
@@ 2 Preliminaries
  • The unit space $\mathcal{G}^{(0)}$ is a Borel subset of $\mathcal{G}$.
  • the source and target maps $s,t:\mathcal{G}\to\mathcal{G}^{(0)}$ are countable-to-one Borel maps,
  • the composition map
    $\{(g,h)\in\mathcal{G}\times\mathcal{G}\mid t(h)=s(g)\}\to\mathcal{G}:(g,h)\mapsto gh$
    and the inversion map $g\in\mathcal{G}\mapsto g^{-1}$ are Borel maps.
See [AD13, Section 2.1] and the references therein for more information about Borel groupoids. A (non-singular) discrete measured groupoid $(\mathcal{G},\mu)$ is a discrete Borel groupoid $\mathcal{G}$ together with a probability measure $\mu$ on $\mathcal{G}^{(0)}$ such that the following two $\sigma$-finite measures on $\mathcal{G}$ are equivalent:
$\mu_{s}(A)$ $=\int_{\mathcal{G}^{(0)}}|\{g\in A\ |\ s(g)=a\}|\ \mathrm{d}\mu(a);$
$\quad\mu_{t}(A)$ $=\int_{\mathcal{G}^{(0)}}|\{g\in A\ |\ t(g)=a\}|\ \mathrm{d}\mu(a).$
@@
Let $(\mathcal{G},\mu)$ be a discrete measured groupoid. To construct the groupoid von Neumann algebra, we start by defining partial isometries $\lambda_{A}\in B(L^{2}(\mathcal{G},\mu_{s}))$ for each bisection $A\subset\mathcal{G}$ by
$\lambda_{A}1_{E}:=1_{A\cdot E}$ (1)
for every subset $E\subset\mathcal{G}$ with $\mu_{s}(E)<\infty$. The groupoid von Neumann algebra is defined by $L(\mathcal{G}):=\{\lambda_{A}\ |\ A\subset\mathcal{G}\text{ Borel bisection}\}^{\prime\prime}$. See [AD13] for more information on the construction. Whenever $A$ and $B$ are bisections, one sees that $\lambda_{A}\lambda_{B}=\lambda_{A\cdot B}$. Note that whenever $A\subset\mathcal{G}$ is a Borel subset with $\mu_{s}(A)<\infty$, it defines an element of $L(\mathcal{G})$ which we also denote by $\lambda_{A}$ as it is defined by the formula (1).
@@
Note that $L(\mathcal{G})$ includes a copy of $L^{\infty}(\mathcal{G}^{(0)},\mu)$ as a regular subalgebra. When $A\subset\mathcal{G}$ is a Borel bisection and $F\in L^{\infty}(\mathcal{G}^{(0)},\mu)$, one has the relation $\lambda_{A}F\lambda_{A}^{*}=(F\circ\sigma_{A}^{-1})\cdot 1_{t(A)}$. Moreover, there is always a faithful normal conditional expectation $E:L(\mathcal{G})\to L^{\infty}(\mathcal{G}^{(0)},\mu)$ given by $a\mapsto a1_{\mathcal{G}^{(0)}}|_{\mathcal{G}^{(0)}}$. The fact that $E(a)\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ follows from Lemma 3.4.
@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
The following proposition is a consequence a theorem of Lusin-Novikov, see e.g. [Kec95, Theorem 18.10], as noted by Anatharaman-Delaroche in [AD13]. We give a proof for convenience.
@@
Since $s:\mathcal{G}\to\mathcal{G}^{(0)}$ is countable-to-one, we can use the Lusin-Novikov Theorem [Kec95, Theorem 18.10] to find a countable collection $(U_{n})_{n}$ of Borel sets $U_{n}\subset\mathcal{G}$ such that $s:U_{n}\to s(U_{n})$ is a Borel isomorphism for every $n$ and such that $\mathcal{G}=\bigsqcup_{n}U_{n}$. Indeed, setting $P:=\{(x,y)\in\mathcal{G}^{(0)}\times\mathcal{G}\ |\ x=s(y)\}$, by Lusin-Novikov there is a partition $P=\bigsqcup_{n}P_{n}$. Denoting by $\pi$ the projection from $\mathcal{G}^{(0)}\times\mathcal{G}$ onto $\mathcal{G}$, set $U_{n}=\pi(P_{n})$. Similarly, there exists a countable collection $(V_{n})_{n}$ of Borel subsets $V_{n}\subset\mathcal{G}$ such that $t:V_{n}\to t(V_{n})$ is a Borel isomorphism for every $n$ and such that $\mathcal{G}=\bigsqcup_{n}V_{n}$. Define =$\mathcal{B}^{\prime}=\{(U_{n}\cap V_{m})\setminus\mathcal{G}_{0}\}_{n,m}$ and set =$\mathcal{B}=\{A\cap B^{-1}\ |\ A,B\in\mathcal{B}^{\prime}\}\cup\{\mathcal{G}^{(0)}\}$. ∎
@@
where $f\cdot 1_{B}$ denotes pointwise multiplication and convergence of the series is in $L^{2}$-norm and where * denotes the convolution product: $f_{1}*f_{2}(g)=\sum_{h:s(h)=s(g)}f_{1}(h)f_{2}(h^{-1}g)$.
@@
$\langle 1_{B},1_{C}\rangle=\langle\lambda_{B}1_{\mathcal{G}^{(0)}},\lambda_{C}1_{\mathcal{G}^{(0)}}\rangle=\langle\lambda_{C^{-1}}\lambda_{B}1_{\mathcal{G}^{(0)}},1_{\mathcal{G}^{(0)}}\rangle=\mu\left(\left(C^{-1}\cdot B\right)\cap\mathcal{G}^{(0)}\right)$
and $\left(C^{-1}\cdot B\right)\cap\mathcal{G}^{(0)}=\emptyset$ whenever $B\neq C$. The decomposition of $f\in L^{2}(\mathcal{G},\mu_{s})$ follows from the fact that $(\mathcal{G},\mu_{s})=\left(\bigsqcup_{B\in\mathcal{B}}B,\sum_{B\in\mathcal{B}}\mu_{s}|_{B}\right)$. Finally, a computation shows that for any $f\in\mathcal{B}$ it holds that
@@
$j:L(\mathcal{G})\to L^{2}(\mathcal{G},\mu_{s}):a\mapsto j(a):=a1_{\mathcal{G}^{(0)}}.$ (3)
It is not difficult to check that $j$ is a continuous linear map. To see that $j$ is injective, denote by $E$ the conditional expectation from $L(\mathcal{G})$ onto $L^{\infty}(\mathcal{G}^{(0)},\mu)$ and define the state $\phi_{\mu}$ on $A(\mathcal{G},\mu)$ by setting
$\phi_{\mu}(\lambda_{A})=\int E(1_{A})\mathrm{d}\mu=\mu(A\cap\mathcal{G}^{(0)})$
whenever $A$ is a Borel subset of $\mathcal{G}$ with $\mu_{s}(A)<\infty$. Then $\phi_{\mu}$ extends to a faithful normal state on $L(G)$ by $\phi_{\mu}(a)=\int E(a)\mathrm{d}\mu$. Now take $a$ and $b$ in $L(\mathcal{G})$ such that $j(a)=j(b)$. Then
$\phi_{\mu}((a-b)^{\ast}(a-b))$ $=\int E((a-b)^{\ast}(a-b))\mathrm{d}\mu$
$=\int E((a-b)^{\ast}(a-b))1_{\mathcal{G}^{(0)}}\mathrm{d}\mu$
$=\int E((a-b)^{\ast}(a-b)1_{\mathcal{G}^{(0)}})\mathrm{d}\mu$
$=\int E((a-b)^{\ast}(j(a)-j(b))\mathrm{d}\mu=0\;.$
Hence the faithfulness of $\phi_{\mu}$ implies that $a-b=0$ so that $j$ is injective.
On $A(\mathcal{G},\mu)$ one sees that $j$ acts as the identity, by which we mean that $j(\lambda_{A})=1_{A}$ for all Borel subsets $A\subset\mathcal{G}$ with $\mu_{s}(A)<\infty$. In fact, it is not difficult to check that $j(\lambda_{A}x\lambda_{B})=\lambda_{A}\rho_{B^{-1}}j(x)$ for all Borel bisections $A,B\subset\mathcal{G}$ and all $x\in L(\mathcal{G})$.
@@
Now suppose $a\in L(\mathcal{G})$ with $\|a\|\leq 1$. Let $(a_{n})_{n}$ be a sequence in $A(\mathcal{G},\mu)$ be a sequence with $\|a_{n}\|\leq 1$ for all $n$ and such that $a_{n}\to a$ in the SOT. By definition, we get that $a_{n}1_{\mathcal{G}^{(0)}}\to a1_{\mathcal{G}^{(0)}}$ in $L^{2}$-norm. Then also $a_{n}1_{\mathcal{G}^{(0)}}\circ s|_{B}^{-1}\to a1_{\mathcal{G}^{(0)}}\circ s|_{B}^{-1}$ in $L^{2}$-norm. Now it suffices to note that from $\|a\|\leq 1$ it follows that $\|a_{n,B}\|_{\infty}=\|a_{n}1_{\mathcal{G}^{(0)}}\circ s|_{B}^{-1}\|_{\infty}\leq 1$. Then, since the $\|\cdot\|_{2}$ and the $\sigma$-strong* topology agree on bounded sets, we get that $a_{n,B}\to a_{B}$ in the SOT, which implies that $a_{B}\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ since all $a_{n,B}\in L^{\infty}(\mathcal{G}^{(0)},\mu)$ by the first part of the proof.
@@
We claim that the net $a_{\mathcal{F}}$ converges to $a$ in the sharp norm. We first show that $\phi_{\mu}((a-a_{\mathcal{F}})^{*}(a-a_{\mathcal{F}}))\to 0$ as $\mathcal{F}$ tends to infinity. This follows from noting that
$\phi_{\mu}((a-a_{\mathcal{F}})^{*}(a-a_{\mathcal{F}}))=\langle j(a-a_{\mathcal{F}}),j(a-a_{\mathcal{F}})\rangle$ $=\left\|\sum_{B\in\mathcal{B}\setminus\mathcal{F}}a^{B}\ast 1_{B}\right\|_{2}$
To see that $\phi_{\mu}((a-a_{\mathcal{F}})(a-a_{\mathcal{F}})^{*}))\to 0$ as $\mathcal{F}$ tends to infinity, note that
@@
so it suffices to check that $(a_{\mathcal{F}})^{*}$ gives the appropriate Fourier decomposition of $a^{*}$. To this end, a simple calculation shows that
@@
Finally, to see that the Fourier coefficients satisfy $E(a\lambda_{B}^{*})=a^{B}$, fix a $B\in\mathcal{B}$ and observe that whenever $\mathcal{F}\subset\mathcal{B}$ is a finite set containing $B$, it holds that $E(a_{\mathcal{F}}\lambda_{B}^{*})=a_{\mathcal{F}}^{B}$. The $E(\cdot\lambda_{B}^{*})$ is a normal map, hence $\|\cdot\|^{\sharp}$-continuous on bounded subsets. Since the net $(a_{\mathcal{F}})$ is bounded in sharp norm by $\|a\|^{\sharp}$, we indeed get that $(a\lambda_{B}^{*})=\lim_{\mathcal{F}}E(a_{\mathcal{F}}\lambda_{B}^{*})=a^{B}$. ∎
@@
Let $\mathcal{G}$ be a discrete Borel groupoid and let $\mathcal{B}$ be a basis for $\mathcal{G}$. Suppose $C\in\mathcal{B}$ is a bisection. The set of conjugates $C\mathcal{B}C^{-1}=\{CBC^{-1}\ |\ B\in\mathcal{B}\}$ also forms a basis for $C\mathcal{G}C^{-1}$. If $\mathcal{B}$ is symmetric, then so is $C\mathcal{B}C^{-1}$.
@@
First let us show that elements of $C\mathcal{B}C^{-1}$ are mutually disjoint. Let $B,D\in\mathcal{B}$. Suppose that $a,b,c,d\in C$ and $x\in B$ and $y\in D$ such that $axb^{-1}=cyd^{-1}$. Note in particular that $t(a)=t(c)$ and that $t(b)=t(d)$. However, each $a,b,c,d$ belong to the same bisection. By injectivity of the target map, we get that $a=c$ and $b=d$. Thus, $x=a^{-1}axb^{-1}b=a^{-1}cyd^{-1}b=c^{-1}cyd^{-1}d=y\in B\cap D$. Therefore, $B=D$. Finally, observe that
@@
We argue by contradiction. Suppose $j(a)$ is not supported on $\operatorname{Iso}(\mathcal{G})$. Then we find a non-null Borel subset $E\subset\mathcal{G}\setminus\operatorname{Iso}(\mathcal{G})$ such that $j(a)(g)\neq 0$ for all $g\in E$. Since $\mathcal{B}$ is countable and all $B\in\mathcal{B}$ are mutually disjoint, there is a $B_{0}\in\mathcal{B}$ such that $\mu_{s}(E\cap B_{0})>0$. Note that for every $x\in s(E\cap B_{0})$ we have that $a_{B_{0}}(x)\neq 0$. By the centrality of $a$, we get for any Borel $F\subset s(E\cap B_{0})$ that $j(\lambda_{F}a)=j(a\lambda_{F})$. Since $F\subset\mathcal{G}^{(0)}$ and $a_{B}\in L^{\infty}(\mathcal{G}^{(0)},\mu)$, a calculation shows
$j(\lambda_{F}a)=\sum_{B\in\mathcal{B}}1_{F\cdot B}\ast a_{B}=\sum_{B}1_{B}\ast(1_{F}\circ\sigma_{B})a_{B}.$
On the other hand, $j(a\lambda_{F})=\sum_{B\in\mathcal{B}}1_{B}\ast a_{B}1_{F}$. Since $a_{B}1_{F}$ and $(1_{F}\circ\sigma_{B})a_{B}$ are both elements of $L^{\infty}(\mathcal{G}^{(0)},\mu)$, looking at the restriction of $j(\lambda_{F}a)$ on each basis element gives us the identity $a_{B}1_{F}=(1_{F}\circ\sigma_{B})a_{B}$ for all $B\in\mathcal{B}$. Hence in particular $1_{F}=(1_{F}\circ\sigma_{B_{0}})$, meaning that for any $x\in s(E\cap B_{0})$, $x\in F$ if and only if $\sigma_{B_{0}}(x)\in F$. Since this holds for any Borel subset $F\subset s(E\cap B_{0})$, we get that $x=t\circ s|_{B_{0}}^{-1}(x)$ for all $x\in s(E\cap B_{0})$. In other words, $E\cap B_{0}\subset\operatorname{Iso}(\mathcal{G})$, which is a contradiction. ∎